a private airplane had been flying for 1 hour when a change of wind direction doubled the plane's average speed. if the entire trip of 384 km took 2.5 hours, how far did the plane travel in the first hour?
Answer by richwmiller(17219)

2.5-1=1.5
x*1+2x(1.5)=384
x+3x=384
4x=384
x=96 kph
in one hour at 96 kph the plane traveled 96 kilometers
59.65 mph
check
96+2(96)81.5=384
96+288=384
384=384
ok